OpenStudy (anonymous):

y '= 1 + 3x - 2y, y (1) = 2 Determine an approximation to y (2) by use of Euler's method with step-by-step length 0.25.

OpenStudy (anonymous):

\[y_{n+1}=y_n+ y'_n(h)\]

OpenStudy (anonymous):

h=1 x=1.25 y(1.25)=y(1)+y'(x,y)(.25) =2 + 2(.25)

OpenStudy (anonymous):

do it three more time

OpenStudy (anonymous):

\[y=(1/4)+(3x/2)+K*1/(e^2x+C)\] Equation itself, the K and C being constants of integration

OpenStudy (anonymous):

This is numerical solution problem